A securities broker advised a client to invest a total of $21,000 in bonds paying 12% interest and in certificates of deposit paying 5 1/4% interest. The annual income from these investiments was $2250. Find out how much was invested at each rate.
Answer by josmiceli(19441)

x = amount invested at 12%
21000 - x = amount invested at 5.25%
.12x + .0525(2100-x) = 2250
1102.50 - .0525x = 2250
.0675x = 1147.5
x = 1147.5/.0675
x = 17000 invested at 12%
21000-x = 4000 invested at 5.25%
check
.12(17000) + .0525(21000-17000) = 2250
2040 + 210 = 2250
2250 = 2250
OK